Methylamine, CH3NH2, has a Kb = 4.40 x 10-4.

What is the pH of a 0.250 M solution of methylamine?
Answer #1

CH3NH2 dissociates as:

CH3NH2 +H2O -----> CH3NH3+ + OH-

0.25 0 0

0.25-x x x

Kb = [CH3NH3+][OH-]/[CH3NH2]

Kb = x*x/(c-x)

Assuming x can be ignored as compared to c

So, above expression becomes

Kb = x*x/(c)

so, x = sqrt (Kb*c)

x = sqrt ((4.4*10^-4)*0.25) = 1.049*10^-2

since x is comparable c, our assumption is not correct

we need to solve this using Quadratic equation

Kb = x*x/(c-x)

4.4*10^-4 = x^2/(0.25-x)

1.1*10^-4 - 4.4*10^-4 *x = x^2

x^2 + 4.4*10^-4 *x-1.1*10^-4 = 0

This is quadratic equation (ax^2+bx+c=0)

a = 1

b = 4.4*10^-4

c = -1.1*10^-4

Roots can be found by

x = {-b + sqrt(b^2-4*a*c)}/2a

x = {-b - sqrt(b^2-4*a*c)}/2a

b^2-4*a*c = 4.402*10^-4

roots are :

x = 1.027*10^-2 and x = -1.071*10^-2

since x can't be negative, the possible value of x is

x = 1.027*10^-2

So, [OH-] = x = 1.027*10^-2 M

use:

pOH = -log [OH-]

= -log (1.027*10^-2)

= 1.9884

use:

PH = 14 - pOH

= 14 - 1.9884

= 12.0116

Answer: 12.01